OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. (2 points) Equation 1 Equation 2 |5x - 6| = -41 |7x + 13| = 27 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (triciaal):

solve each eqn eqn 1 -(5 x - 6) = -41 -5 x = -41 -30 =-71 x = 71/5 (5 x - 6)= -41 5 x = -41 + 30 = -11 x = -11/5 eqn 2 |7x + 13| = 27 -7x -13 = 27 7 x = 27 + 13 x = -40/7 7 x + 13 = 27 x = 27 - 13 x = 14/7 = 2

OpenStudy (anonymous):

so which one one these is it ? Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.